# ELemental abundances of Planets and brown dwarfs Imaged around Stars (ELPIS): II. The Jupiter-like Inhomogeneous Atmosphere of the First Directly Imaged Planetary-Mass Companion 2MASS 1207 b

### v1

@@ 2. $\bullet$
    An expanded grid of chemical equilibrium abundances for petitRADTRANS, which extends the temperature range and incorporates additional atomic and molecular species (Section 5.2.2).
@@ 3. $\bullet$
Future observational studies of 2MASS 1207 b, such as JWST MIRI spectroscopy (GTO program 1270) and time-series NIRSpec spectroscopy (GO program 3181), will provide new insights into this remarkable planetary-mass companion. Additionally, detailed atmospheric studies of its host, 2MASS 1207 A, will offer a crucial context for understanding formation pathways and the evolutionary history of this planetary system.
We thank the anonymous referee for suggestions that improved the manuscript. Z.Z. thanks Gaël Chauvin, Elena Manjavacas, Andy Skemer, and Xi Zhang for helpful discussions during the early stages of this work. Z.Z. acknowledges the support of the NASA Hubble Fellowship grant HST-HF2-51522.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. We acknowledge the use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. The data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/gp39-v372 (catalog DOI: 10.17909/gp39-v372). Z.Z. warmly acknowledges the love and support of his wife, Yanxia Li, and the arrival of our daughter, Allison Muzi Zhang, whose presence has brought immense joy and inspiration during the completion of this work.

### v2

@@ 2. $\bullet$
    An expanded grid of chemical equilibrium abundances [^10] https://doi.org/10.5281/zenodo.15654694 for petitRADTRANS, which extends the temperature range and incorporates additional atomic and molecular species (Section 5.2.2).
@@ 3. $\bullet$
Future observational studies of 2MASS 1207 b, such as JWST MIRI spectroscopy (GTO program 1270) and time-series NIRSpec spectroscopy (GO program 3181), will provide new insights into this remarkable planetary-mass companion. In particular, follow-up atmospheric retrieval analyses of combined NIRSpec and MIRI data for 2MASS 1207 b should account for both its inhomogeneous atmosphere and the possible presence of its circumsecondary disk, which, if present, could contribute to the flux at MIRI wavelengths. Additionally, detailed atmospheric studies of its host, 2MASS 1207 A, will offer a crucial context for understanding formation pathways and the evolutionary history of this planetary system.
We thank the anonymous referee for suggestions that improved the manuscript. Z.Z. thanks Gaël Chauvin, Gabriel-Dominique Marleau, and Elena Manjavacas for helpful comments and thanks Andy Skemer and Xi Zhang for discussions during the early stages of this work. Z.Z. acknowledges the support of the NASA Hubble Fellowship grant HST-HF2-51522.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. We acknowledge the use of the lux supercomputer at UC Santa Cruz, funded by NSF MRI grant AST 1828315. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular, the institutions participating in the Gaia Multilateral Agreement. The data presented in this article were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via https://doi.org/10.17909/jzze-5611 (catalog DOI: 10.17909/jzze-5611). Z.Z. warmly acknowledges the love and support of his wife, Yanxia Li, and the arrival of our daughter, Allison Muzi Zhang, whose presence has brought immense joy and inspiration during the completion of this work.